# Understanding Epistemic Language with a Bayesian Theory of Mind

### v1

@@ Abstract
How do people understand and evaluate claims about others’ beliefs, even though these beliefs cannot be directly observed? In this paper, we introduce a cognitive model of epistemic language interpretation, grounded in Bayesian inferences about other agents’ goals, beliefs, and intentions: a language-augmented Bayesian theory-of-mind (LaBToM). By translating natural language into an epistemic “language-of-thought”, then evaluating these translations against the inferences produced by inverting a probabilistic generative model of rational action and perception, LaBToM captures graded plausibility judgments about epistemic claims. We validate our model in an experiment where participants watch an agent navigate a maze to find keys hidden in boxes needed to reach their goal, then rate sentences about the agent’s beliefs. In contrast with multimodal LLMs (GPT-4o, Gemini Pro) and ablated models, our model correlates highly with human judgments for a wide range of expressions, including modal language, uncertainty expressions, knowledge claims, likelihood comparisons, and attributions of false belief.
@@
Figure 1: Overview of our model, a Language-augmented Bayesian Theory of Mind (LaBToM). In this example, our model evaluates the plausibility of epistemic language about a player trying to find keys in boxes so as to reach one of four valuable gems. To do so, (a) we use an LLM to translate natural language into an unambiguous epistemic language of thought (ELoT). (b) We then use our Bayesian theory-of-mind (BToM) module to infer the agent’s goal $g$ and probabilistic beliefs $b$, along with the true environment state $s$. This allows us to evaluate the normalized likelihood $\bar{L}$ (see Section 3.2.3) of an epistemic statement about the agent’s present or past beliefs.
@@ 1 Introduction
People regularly use and interpret language about other agents’ beliefs, evaluating rich linguistic constructions that may involve claims about what others’ consider necessary (“Grace thinks that if Katie didn’t eat the cookie, it must have been Jane”), what they find more probable (“Tom believes Sam is more likely to win the election than the others”), or the relationship of their beliefs to the world (“John didn’t know that today was a holiday”). But given that these beliefs are not directly observable, how do people evaluate the truth or plausibility of epistemic claims? Philosophers and linguists have long investigated the semantics of epistemic language Hintikka (1962); Partee (1973); Loar (1981), offering detailed theories of how epistemic claims relate to the sets of worlds deemed possible by an agent, and how the truth of these claims derive from their constituent predicates or propositions Von Fintel and Heim (2011). However, these theories do not explain how people understand and evaluate epistemic language in its context of utterance. If someone says “Alice believes it might rain”, how does her behavior (e.g. bringing an umbrella) render that statement more or less plausible?
@@
In this paper, we introduce a cognitive model of how humans interpret epistemic language in context (Figure 1), grounded in inferences about what agents believe given their actions and observations. We build upon a framework known as Bayesian Theory-of-Mind (BToM), which casts human mentalizing as Bayesian inference over a generative model of rational action and perception Baker et al. (2017); Jara-Ettinger et al. (2019). We combine this framework with the compositionality afforded by probabilistic extensions of the language-of-thought hypothesis (Piantadosi, 2011; Goodman and Lassiter, 2015), developing an epistemic language of thought (ELoT) to represent how others represent the world. Using large language models (LLMs) as flexible semantic parsers, we translate natural language into this ELoT representation (Wong et al., 2023), allowing us to quantitatively evaluate epistemic claims against BToM inferences. This gives us our full model: A language-augmented Bayesian theory-of-mind (LaBToM).
@@
To evaluate our model, we run an experiment where participants watch animations of a player solving a gridworld puzzle called Doors, Keys, & Gems Zhi-Xuan et al. (2020). In these puzzles (Figure 1b), the player has to pick up (single-use) keys that may be hidden in boxes, using them to unlock doors of the same color to reach one of four valuable gems. The player’s beliefs and goals are unknown to our participants, so they must infer these mental states. We ask one set of participants to write sentences describing the past and present beliefs of the player, collecting a rich dataset of epistemic language that includes modal verbs, uncertainty expressions, knowledge claims, and descriptions of false beliefs. We then task another set of participants with evaluating these statements, asking them to rate how likely a statement is given their observations. We find that the inferences produced by our LaBToM model correlate highly with these ratings. In contrast, ablated BToM models and multimodal LLM baselines fail to explain human judgments, illustrating the importance of a coherent theory-of-mind for understanding epistemic language.
@@ 2 Related Work
Semantics of Epistemic Language.
Sentences about belief and knowledge have been studied at length in philosophy and linguistics Hawthorne et al. (2016); Egan and Weatherson (2011); Yalcin (2007). Standard treatments associate belief sentences with sets of compatible worlds Hintikka (1962); Von Fintel and Heim (2011), while more recent work grounds the gradedness of belief claims in probabilities Lassiter (2017); Moss (2015). While our model draws upon these frameworks, such theories do not explain how people evaluate belief claims in context.
Cognitive Models of Language Interpretation.
In contrast to the above line of research, our goal is to build a cognitive model of epistemic language understanding, drawing upon accounts of word meaning that prioritize the role of mental states Loar (1981); Block (1987); Lake and Murphy (2023), including cognitive semantics Lakoff (1988); Jackendoff (2003), functional role semantics Harman (1982), and work that grounds linguistic meaning in a (probabilistic) language of thought (Fodor et al., 1975; Goodman and Lassiter, 2015; Wong et al., 2023; Zhang et al., 2023).
@@
To model how epistemic sentences relates to (inferred) mental states, our paper builds upon work in Bayesian Theory-of-Mind Baker et al. (2017); Zhi-Xuan et al. (2022), along with related work on epistemic action understanding Croom et al. (2023); Shvo et al. (2020), leveraging the connection between symbolic representations of the world (used in planning and inverse planning McDermott et al. (1998)) and our representations of others’ minds.
@@
With recent advances in the capabilities of neural large language models (LLMs), some researchers have suggested that LLMs might serve as cognitive models Binz and Schulz (2023), including as models of theory-of-mind Strachan et al. (2024). However, while LLMs achieve human-like performance on some ToM tasks, they do not reliably generalize to simple alterations Shapira et al. (2024), multi-step conversations Kim et al. (2023), or multi-modal scenarios Jin et al. (2023); Ying et al. (2024a). Our model represents an alternative structured approach, which leverages LLMs as flexible translators between natural language and formal representations of meaning Wong et al. (2023); Ying et al. (2023a, b); Zhi-Xuan et al. (2024b).
@@ 3 Computational Model
Our LaBToM model comprises two interlinked modules for reasoning about beliefs. The first module (Figure 1a), an epistemic language of thought, models our capacity to compositionally represent concepts that describe the world (including the contents of other agents’ minds) by combining more basic concepts into richer thoughts and expressions Goodman and Lassiter (2015), and how we translate such thoughts from natural language Wong et al. (2023). The second module (Figure 1b), a Bayesian theory-of-mind Baker et al. (2017), captures our intuitive inferences about others’ mental states via probabilistic inference over a generative model of how agents update their beliefs and act towards their goals. By combining these two modules, epistemic language understanding can be modeled by first translating natural language into our ELoT representation, then evaluating these representations against inferences produced by rational mentalizing.
To represent epistemic sentences in a way that admits systematic compositionality, we introduce a formal language (Table 1) that serves as our epistemic language of thought. This language can be viewed as a form of epistemic logic Gochet and Gribomont (2006) with a richer set of modal operators. We adopt a degree-based semantics for epistemic modals grounded in probability comparisons, due to its advantages in handling the gradability of epistemic expressions Lassiter (2010); Moss (2015), but crucially also because it allows us to evaluate formulas using probabilities inferred by our BToM module.
@@ 3.1.1 Representing Epistemic Formulas
On top of this base language, we introduce a set of epistemic expressions to model assertions of belief, knowledge, or modal qualifications thereof (Table 1a). The semantics of these expressions are ultimately grounded in the probability function $\textbf{Pr}(A,\phi)$ — the probability assigned by agent $A$ to formula $\phi$ — and comparisons against term-specific thresholds (Table 1b). For example, the operator $\textsf{might}(\phi)$ takes a formula $\phi$ and returns a function $\lambda A.\textbf{Pr}(A,\phi)\geq\theta_{\textsf{might}}$. Combined with the operator $\textsf{believes}_{\textsf{modal}}(A,F)=F(A)$, we can express the claim that $A$ believes it might be that $\phi$ as $\textsf{believes}_{\textsf{modal}}(A,\textsf{might}(\phi))$. This term can be lowered into the probability comparison $\textbf{Pr}(A,\phi)\geq\theta_{\textsf{might}}$, which uses $\theta_{\textsf{might}}$ as the threshold instead of $\theta_{\textsf{believes}}$ due to the way our operators compose. Similar to CCG-based semantic frameworks, our ELoT module also supports comparatives Haruta et al. (2019) and typed higher-order terms Mineshima et al. (2015); Martínez-Gómez et al. (2016).
@@ 3.1.2 Translating Epistemic Language
Epistemic formulas $\varphi$ in our ELoT representation are unambiguous and precise. In contrast, natural language might be used to communicate the same epistemic information in multiple ways. To handle this diversity of surface forms, we make use of LLMs as general-purpose semantic parsers (Shin et al., 2021) that translate a wide variety of expressions into formal representations, substituting for the human ability to flexibly interpret natural language in-context Wong et al. (2023). In particular, we prompt an LLM (Gemini Flash) with example translations from English to our ELoT representation, and append a new sentence $\sigma$ to translate. We then sample up to $n_{\sigma}$ outputs, rejecting any which are syntactically or semantically invalid, retaining the first formula $\varphi$ that is successfully translated. (See Appendix for the translation prompt and other details.)
@@ 3.2 Inferring and Evaluating Beliefs with a Bayesian Theory-of-Mind
Our ELoT module gives us a way to interpret belief claims in terms of the probability $\textbf{Pr}(A,\phi)$ an agent $A$ assigns to sentences $\phi$. But how is the value of $\textbf{Pr}(A,\phi)$ computed? This is the function of our BToM module: By modeling the functional role that belief plays in guiding an agent’s actions, along with the causal influence of the agent’s perceptions on their beliefs, an observer can infer what the agent thinks based on what the agent sees and does. Following the structure of Partially Observable Markov Decision Processes (POMDPs) Kaelbling et al. (1998), this theory of approximately rational agency can be formalized as a probabilistic generative model:
@@ 3.2.2 Joint Inference of Goals and Beliefs
Computing this posterior is intractable in general due to the large space of possible initial beliefs $b_{0}$, which may in turn be defined over a large space of environment states $s_{0}$. However, if the space of possible beliefs and states is sufficiently small, it can be reasonable to model human observers as approximating the Bayesian ideal Blokpoel et al. (2010). Therefore, we consider only the set of initial states $\mathcal{S}_{0}$ consistent with the initial observation $o_{0}$, and a discrete set $\mathcal{B}_{0}$ of possible beliefs $b_{0}$ sufficient to model comparative likelihood claims (e.g. “The key is more likely in box 1 than 2.”). Specifically, we consider all beliefs formed by distributing $k$ particles across $n_{s}:=|\mathcal{S}_{0}|$ states, resulting in $n_{b}:=|\mathcal{B}_{0}|=\binom{n_{s}+k-1}{k}$ distributions. We then perform exact Bayesian inference over all combinations of goals $g$, initial beliefs $b_{0}$, and states $s_{0}$, which we implement as a variant of Sequential Inverse Plan Search (Zhi-Xuan et al., 2020) using the Gen probabilistic programming system (Cusumano-Towner et al., 2019). More algorithmic details are provided in the Appendix.

### v2

# Understanding Epistemic Language with a Language-augmented Bayesian Theory of Mind

@@ Abstract
How do people understand and evaluate claims about others’ beliefs, even though these beliefs cannot be directly observed? In this paper, we introduce a cognitive model of epistemic language interpretation, grounded in Bayesian inferences about other agents’ goals, beliefs, and intentions: a language-augmented Bayesian theory-of-mind (LaBToM). By translating natural language into an epistemic “language-of-thought” with grammar-constrained LLM decoding, then evaluating these translations against the inferences produced by inverting a generative model of rational action and perception, LaBToM captures graded plausibility judgments of epistemic claims. We validate our model in an experiment where participants watch an agent navigate a maze to find keys hidden in boxes needed to reach their goal, then rate sentences about the agent’s beliefs. In contrast with multimodal LLMs (GPT-4o, Gemini Pro) and ablated models, our model correlates highly with human judgments for a wide range of expressions, including modal language, uncertainty expressions, knowledge claims, likelihood comparisons, and attributions of false belief.
@@
Figure 1: Overview of our model, a Language-augmented Bayesian Theory of Mind (LaBToM). Here, our model evaluates the plausibility of epistemic language about a player character trying to find keys in boxes so as to reach one of four goals (gems with different shapes/colors, $\textsf{g}_{1}$$\textsf{g}_{4}$) which may be locked behind doors. (Top) We translate natural language statements about the player’s initial ($\sigma_{1}$) and current beliefs ($\sigma_{2}$) into an unambiguous epistemic language of thought (ELoT) via grammar-constrained LLM parsing. ELoT statements (i.e. epistemic formulas $\varphi$) are interpreted with a probability-based semantics (lowered formulas $\varphi^{\prime}$). (Bottom) We use our Bayesian theory-of-mind (BToM) module to produce inferences (bar charts) about the environment state (top left), the player’s goal (bottom left), and the player’s belief state (right) at each step $t$, given observations across time. Each possible belief state $\textsf{b}_{i}$ is itself a distribution over environment states $\textsf{s}_{j}$, where state $\textsf{s}_{j}$ corresponds to a blue key being in box $j$. (Middle) We evaluate the posterior probability of each ELoT statement $\varphi$ from the BToM inferences at each step $t$ (under a 50-50 prior over statement truth). Statement $\sigma_{1}$ (“The player thought that the blue key must be in box 3.”) increases in probability from $t$$=$$6$ to $t$$=$$9$, then stays high, since it becomes clear from the player looking in box 3 at $t$$=$$9$ that they initially thought a key must be in box 3. Statement $\sigma_{2}$ (“The player believes that a blue key is more likely to be in box 1 than box 2.”) decreases in probability as the player walks away from box 1 at $t$$=$$6$ through $t$$=$$9$. However, when the player finds box 3 empty, then moves past box 2 up to box 1 at $t$$=$$15$, it becomes much more probable that the player currently thinks a key is more likely to be in box 1 rather than box 2.
@@ 1 Introduction
People regularly use and interpret language about other agents’ beliefs, evaluating rich linguistic constructions that may involve claims about what others’ consider possible (“Grace thinks Katie might have eaten the cookie”), what they find more probable (“Tom thinks Sam is the most likely to win the election”), or the relationship of their beliefs to the world (“John didn’t know that today was a holiday”). But since these beliefs are not directly observable, how do people judge the truth or acceptability of epistemic claims? Philosophers and linguists have long investigated the semantics of epistemic language Hintikka (1962); Partee (1973); Loar (1981), offering compositional theories of how epistemic claims relate to the sets of worlds deemed possible by an agent Von Fintel and Heim (2011). However, these theories do not explain how people ground epistemic language in its context of utterance. If someone says “Alice believes it might rain”, how does her behavior (e.g. bringing an umbrella) render that statement more or less plausible or acceptable?
@@
In this paper, we present a computational model of how people interpret epistemic language in context, grounded in inferences about what other agents believe (Figure 1) — a Language-augmented Bayesian theory-of-mind (LaBToM). We build upon Bayesian theory-of-mind (BToM), a framework which casts human mentalizing as probabilistic inference over a generative model of rational action and perception Baker et al. (2017); Jara-Ettinger et al. (2020). We combine this framework with the compositionality afforded by probabilistic extensions of the language-of-thought hypothesis Piantadosi (2011); Goodman and Lassiter (2015), developing an epistemic language of thought (ELoT) to represent how others represent the world. We parse natural language to this representation using grammar-constrained sequential Monte Carlo decoding Loula et al. (2025) of large language models (LLMs), defining a flexible but structured probabilistic mapping from epistemic language to epistemic concepts Wong et al. (2023). This allows us to quantify the plausibility of epistemic sentences against BToM inferences.
@@
To evaluate our model, we run an experiment where participants watch animations of a player solving a gridworld puzzle called Doors, Keys, & Gems Zhi-Xuan et al. (2020). In these puzzles, the player has to pick up (single-use) keys that may be hidden in boxes, using them to unlock doors of the same color to reach one of four valuable gems. The player’s beliefs and goals are unknown to our participants, so they must infer these mental states. We ask one set of participants to write sentences describing the past and present beliefs of the player, collecting a rich dataset of epistemic language that includes modal verbs, uncertainty expressions, knowledge claims, and descriptions of false beliefs. We task another set of participants with evaluating these statements, asking them to rate how likely a statement is given the context. We find that the inferences produced by our LaBToM model correlate highly with these human ratings. In contrast, using a lower-level formalization of epistemic claims leads to substantially more parsing errors than our ELoT representation, and ablated BToM models and multimodal LLM baselines fail to explain human sentence judgments. These findings illustrate the importance of a conceptual vocabulary that is aligned with the structure of epistemic language, and the need for a coherent theory-of-mind to systematically interpret such language.
@@ 2 Related Work
Model-theoretic formal semantics of epistemic language.
Sentences about what agents believe or know have been studied extensively across philosophy and linguistics. Both epistemic modals (Yalcin, 2007; Coates, 1987; Egan and Weatherson, 2011; MacFarlane, 2011) and propositional attitudes (Frege, 1948; Russell, 1903; Schiffer, 2003) complicate compositional and model-theoretic approaches to assigning truth conditions. Standard treatments associate belief sentences with sets of compatible worlds Hintikka (1962); Von Fintel and Heim (2011), though the role of an agent’s mental states in a truth-conditional semantics remains debated (Egan and Weatherson, 2011). More recent work grounds the gradedness of belief claims in probability Lassiter (2017); Moss (2015), which we follow in this paper. Unlike purely model-theoretic approaches, however, our model relates epistemic sentences to a functional theory of agents’ mental states (Loar, 1981), explaining how people ground these sentences in agent behavior.
Cognitive approaches to (epistemic) language interpretation.
Our model draws on accounts of propositional attitudes (A. Fodor, 1981) and word meaning that emphasize the role of mental states Loar (1981); Block (1987); Lake and Murphy (2023) and the interface between language and conceptual content, as in cognitive or psychosemantics Fodor (1987); Lakoff (1988); Jackendoff (2003) and functional role semantics Harman (1982). We build most closely upon theories that ground linguistic meaning in a (probabilistic) language of thought (Fodor et al., 1975; Goodman and Lassiter, 2015; Wong et al., 2023; Zhang et al., 2023), and models that map sentences to grounded symbolic representations of agents’ behavior (Artzi and Zettlemoyer, 2013).
@@
To relate epistemic sentences to (inferred) mental states, we build on the Bayesian Theory-of-Mind framework Baker et al. (2017); Zhi-Xuan et al. (2022), and related work on epistemic action understanding Croom et al. (2023); Shvo et al. (2020). BToM provides a functional theory of propositional attitudes like knowledge and belief, leveraging the connection between symbolic representations of the world (used in (inverse) planning McDermott et al. (1998)) and our representations of others’ minds.
@@
With recent advances in LLM capabilities, some researchers have suggested that LLMs might serve as cognitive models Binz and Schulz (2023), including as models of theory-of-mind Strachan et al. (2024), thereby implicitly capturing the semantics of belief sentences Piantadosi and Hill (2022). However, while LLMs perform well on some ToM tasks, they do not reliably generalize Shapira et al. (2024) to richer multi-step Kim et al. (2023) or multi-modal contexts Jin et al. (2023); Ying et al. (2024a). Our model instead uses (grammar-constrained) LLMs as a flexible mappings between language and formal meaning representations Wong et al. (2023); Ying et al. (2023a, b); Zhi-Xuan et al. (2024b), tying language to an explicit semantics of epistemic concepts.
@@ 3 Computational Model
Our LaBToM model comprises two interlinked modules. The first module (Figure 1, Top), an epistemic language of thought (ELoT), models our capacity to compositionally represent the world (including the contents of others’ minds) by combining more basic concepts into richer thoughts and expressions Goodman and Lassiter (2015), and how we flexibly translate such thoughts from natural language Wong et al. (2023). The second module (Figure 1, Bottom), a Bayesian theory-of-mind Baker et al. (2017), captures our intuitive inferences about others’ minds via Bayesian inference over a generative model of how agents update their beliefs and act towards their goals. Epistemic language understanding can thus be modeled by mapping language into ELoT formulas, which we evaluate against the inferences produced by rational mentalizing.
@@ 3.1 Interpreting Belief Sentences with an Epistemic Language of Thought
To represent epistemic concepts in a way that mirrors the structure of natural language, we introduce a formal language (Table 1) as our epistemic language of thought. Drawing upon the formal semantics of epistemic language Hintikka (1962); Lassiter (2010), we adopt a compositional degree-based semantics for epistemic concepts grounded in probability comparisons Moss (2015); Lassiter (2017). This allows us to quantitatively evaluate an epistemic expression using probabilities inferred by our BToM module.
@@ 3.1.1 Representing Epistemic Formulas
On top of this base language $\Phi$, we introduce epistemic expressions $\varphi\in\mathcal{E}$ to model assertions of belief, knowledge, or modal qualifications (Table 1a). Following Lassiter (2017), the semantics of these expressions are grounded in the probability function $\textbf{Pr}(A,\phi)$ — the probability assigned by agent $A$ to formula $\phi$ — and comparisons with term-specific thresholds (Table 1b). For example, the operator $\textsf{might}(\phi)$ takes a first-order formula $\phi\in\Phi$, returning a function $\lambda A.\textbf{Pr}(A,\phi)\geq\theta_{\textsf{might}}$. Combined with the operator $\textsf{believes}_{\textsf{modal}}(A,F)=F(A)$, we can express the claim that $A$ believes it might be that $\phi$ as $\varphi$ $=$ $\textsf{believes}_{\textsf{modal}}(A,\textsf{might}(\phi))$. This formula $\varphi$ can be lowered to the probability comparison $\textbf{Pr}(A,\phi)\geq\theta_{\textsf{might}}$, which uses $\theta_{\textsf{might}}$ as the threshold instead of $\theta_{\textsf{believes}}$ due to how our composition rules. We call this latter representation a lowered formula.
A key aspect of ELoT expressions is that they represent epistemic concepts at a similar level of granularity as our tested language (English), simplifying the mapping to ELoT propositions. In contrast, our lowered representation is just as expressive, but fails to match the structure of natural epistemic language. In our experiments, we show that translating into ELoT significantly improves the accuracy of semantic parsing.
@@ 3.1.2 Translating Epistemic Language
Natural language is varied and imprecise, with possibly many sentences $\sigma$ mapping to the same ELoT formula $\varphi$ (and vice versa). To handle this diversity, we perform semantic parsing via grammar-constrained LLM decoding Shin et al. (2021); Scholak et al. (2021); Willard and Louf (2023), producing a flexible context-sensitive mapping between natural language and ELoT expressions Wong et al. (2023). Specifically, we use sequential Monte Carlo (SMC)-based grammar-constrained sampling (Loula et al., 2025), since it avoids the failure modes of beam search and greedy token-masking Lew et al. (2023b). We prompt an LLM (LLaMa 3.1 8B) with example translations $\mathcal{D}$ from English to ELoT and a sentence $\sigma$ to translate, using SMC to approximate the distribution over completions constrained to the ELoT grammar $G_{\mathcal{E}}$:
$\varphi$ $\sim P_{\text{LLM}}(\varphi|\sigma,\mathcal{D},\varphi\in G_{\mathcal{E}})$ (1)
We run SMC with $n_{\sigma}$ samples. This produces weighted samples $\{(\varphi_{i},w_{i})\}_{i=1}^{n_{\sigma}}$, and we use the top-weighted sample as our ELoT translation $\varphi$ [^1] The full set of SMC samples can instead be used to handle ambiguous sentences with multiple plausible ELoT translations. We leave the study of this to future work.
@@ 3.2 Inferring and Evaluating Beliefs with a Bayesian Theory-of-Mind
Our ELoT semantics reduces belief expressions to the probability $\textbf{Pr}(A,\phi)$ an agent $A$ assigns to a sentence $\phi$. But where does $\textbf{Pr}(A,\phi)$ come from? This is the function of our BToM module: By modeling the functional role that belief plays in guiding actions, along with the influence of perceptions on beliefs, an observer can infer what the agent thinks based on what the agent sees and does. Following the structure of Partially Observable Markov Decision Processes (POMDPs) Kaelbling et al. (1998), this theory of approximately rational agency can be formalized as a probabilistic generative model:
@@ 3.2.2 Joint Inference of Goals and Beliefs
To ensure tractable posterior inference, we considered only the set of initial states $\mathcal{S}_{0}$ consistent with the initial observation $o_{0}$, and a discrete set $\mathcal{B}_{0}$ of possible beliefs $b_{0}$ sufficient to model comparative likelihood claims (e.g. “The key is more likely in box 1 than 2.”). Specifically, we consider all beliefs formed by distributing $k$ particles across $n_{s}:=|\mathcal{S}_{0}|$ states, resulting in $n_{b}:=|\mathcal{B}_{0}|=\binom{n_{s}+k-1}{k}$ distributions. We then perform exact Bayesian inference over all combinations of goals $g$, initial beliefs $b_{0}$, and states $s_{0}$, which we implement as a variant of Sequential Inverse Plan Search Zhi-Xuan et al. (2020) using the Gen probabilistic programming system Cusumano-Towner et al. (2019). More algorithmic details are provided in the Appendix.
3.2.3 Evaluating Epistemic Sentences
By inferring the agent’s belief history $b_{0:T}$, we can compute the probability $\mathbf{Pr}(A,\phi)$ of a formula $\phi$ at time $t$ as the expected truth value under $b_{t}$:
However, observers do not have access to the true $s_{t}$ or $b_{t}$, only inferences about them. As such, we model human evaluation of a sentence $\varphi$ as a probabilistic judgment given their inferences:
4 Experiments
To evaluate our model on a diverse dataset of epistemic language (Table 2), we conducted a two-part human experiment. We first recruited participants to write English sentences describing the current and initial beliefs of a player character as it navigated a gridworld puzzle that required finding keys hidden in boxes. Next, we asked two groups of participants to rate how likely these sentences were to be true given the player’s behavior, with one group rating sentences about the player’s current beliefs, and the other rating sentences about initial beliefs.
With this data, we evaluated our model by: (i) assessing the translation accuracy of our ELoT module, investigating the importance of our ELoT representation (vs. the lowered form) and its impact on grammar-constrained LLM decoding; (ii) testing our full LaBToM model in its ability to capture human interpretation of epistemic language, comparing human ratings against LaBToM inferences. Model and experiment code is available at https://osf.io/xq9c4/.
4.1 Scenario Construction